Tech Meeting Minutes 20100923
Present: Graeme, Sam, Wahid, Mike, Mark, David
|Table of contents|
- Patched WNs, main user UI, and CREAM CE last week.
- Will patch 'specialist' UIs today.
- Mike is rolling a rebuild across the WNs in batches of 10.
- Sam will patch SL5 disk servers and DPM headnode. This will require a very short downtime.
- Need to patch site BDII and rebuild CREAM box.
- Systems team are rolling SL kernel across the WNs.
- Have patched grid/public systems.
- ACTION: Need to do a syncat dump of the DPM to clean dark data.
- Failing top level BDII. No one uses this and the BDII infrastructure is much better now, so we decided to decommission it. ACTION: Mike.
- ATLAS testing is going well. Watch out for 999 bug: http://grid.pd.infn.it/cream/field.php?n=Main.KnownIssues, bug 70287.
- Glasgow will convert 1 lcg-CE to CREAM. ACTION: Mark.
- ECDF have a CREAM box, currently disabled waiting kernel patch.
- Durham ?
- Low priority action to move from glite-MON to glite-APEL by end of year.
- Glasgow: DC will look at this next week. ACTION: David.
- ECDF: Andy will do it post-CHEP. ACTION: Andy.
- Durham: ? ACTION.
- Clustervision have handed over disk servers to Glasgow team. David and Sam will burn them in.
- ECDF, Wahid has new disk servers. Will be integrated into current DPM, but pre-CHEP Wahid will use them for various tests.
- Durham: Tender news?
- Next meeting Thursday 30th, 1100. http://indico.cern.ch/conferenceDisplay.py?confId=108379.